OpenStudy (anonymous):

so the x-intercept of 5y=x+2 is

OpenStudy (anonymous):

x-intercept is where y = 0 5(0) = x +2 0 = x + 2 -2 = x

OpenStudy (anonymous):

5y=x+2 To find the x-intercept, substitute in 0 for y and solve for x. 5(0)=x+2 Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. x+2=5(0) Multiply 5 by the 0 inside the parentheses. x+2=5*0 Multiply 5 by 0 to get 0. x+2=0 Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. x=-2 To find the y-intercept, substitute in 0 for x and solve for y. 5y=(0)+2 Remove the parentheses around the expression 0. 5y=0+2 Adding 0 to an expression does not change the value of the expression. 5y=2 Divide each term in the equation by 5. (5y)/(5)=(2)/(5) Cancel the common factor of 5 in (5y)/(5). (<X>5<x>y)/(<X>5<x>)=(2)/(5) Remove the common factors that were cancelled out. y=(2)/(5) These are the x and y intercepts of the equation 5y=x+2. x=-2, y=(2)/(5)

OpenStudy (anonymous):

so: Intercepts x=-2 y=(2)/(5)

OpenStudy (anonymous):

yes.